OpenStudy (anonymous):

How to simplify Radical8

OpenStudy (anonymous):

\[\sqrt8=\sqrt{4 * 2} = \sqrt{4}*\sqrt{2}=2 \sqrt2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Woah how did you get that

OpenStudy (anonymous):

 I see thank you